A firm uses a serial assembly system and needs answers to the following:

a. An output of 900 units per shift (10.50 hours) is desired for a new processing system. The system requires product to pass through four stations where the work content at each station is 30 seconds. What is the required cycle time for such a system? (Round your answer to the nearest whole number.)

b. How efficient is your system with the cycle time you calculated? (Round your answer to 1 decimal place.)

c. Station 3 changes and now requires 50 seconds to complete. What will need to be done to meet demand (assume only 10.50 hours are available)? What is the efficiency of the new system? (Round your answer to 1 decimal place.)
Answer #1

Answer a:

desired output = 900 units per 10.5 hours=  10.5*60*60/900= 42 secs per unit

Hence the workstation cycle time required is 42 secs

Answer b: Total task time= 30 secs*4 = 120 secs, No. of workstation= 4

Efficiency= total task time / no. of workstation* cycle time = 120/(4*42)= 0.7142 or 71.4%

Answer c: If station 3 now takes 50 secs, than the cycle time will be 50 secs. We need to employ 2 parallel workstations at station 3 so that the effective cycle time becomes its half that is 50/2= 25 secs.

If station 3 takes 50 secs, total task time= 30+30+50+30= 140

cycle time= 50 secs

Efficiency= total task time / no. of workstation* cycle time = 140/(4*50)= 0.7 or 70.0%